OpenStudy (anonymous):

7/19+1/2=7/19+2/1=9/19 would this be correct to lowest terms?

OpenStudy (anonymous):

...lowest terms means? ... do you have to solve it or? .. I don't understand English well enough . :(

Directrix (directrix):

7/19 + 1/2 = 7/19 +2/1 = 9/19 --> Let's see. ==================== 7/19 + 1/2 are not like terms. One has a last name of 19 and the other 2. A common denominator would be 19*2 or 38. So, 7/19 + 1/2 = 7/19 (2/2) + (1/2) (19/19) Note that 2/2 and 19/19 are both equal to 1 so no values are changed. 7/19 + 1/2 = 14 / 38 + 19 / 38. Now we have like terms and can add. 7/19 + 1/2 = 14 / 38 + 19 / 38 = 33 /38 --> Answer, please check.

OpenStudy (anonymous):

@Directrix I get the same too...

OpenStudy (anonymous):

\[\Large \frac{7}{19}+\frac12=\frac{7\cdot 2}{19\cdot 2}+\frac{1\cdot 19}{2\cdot 19}\] \[\Large \frac{14}{38}+\frac{ 19}{38}=\frac{14+19}{38}=\frac{33}{38} \]

OpenStudy (anonymous):

Thank you guys

OpenStudy (anonymous):

you're welcome...